OpenStudy (anonymous):

Okay so for exponential and logarithmic equations I'm totally stuck on how to solve for x on this one. 23^x = 15

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=23%5Ex+%3D+15+solve+for+x&dataset= Does this help you?

OpenStudy (anonymous):

Hmm I think it's on the right path but I'm not exactly sure what it's saying. I probably should have given the options for answers. A. 1.565 B. 2.678 C. 3.187 D. 1.302

OpenStudy (anonymous):

I'm sorry I've been working on this for half an hour so my brain's a little fried.

OpenStudy (anonymous):

That's okay! None of those answers will give you 15. X will need to roughly equal 0.86 in order to bring you to 15.

OpenStudy (wolf1728):

23^x = 15 taking logs of both sides x * log 23 = log 15 x = log 15 / log 23 enough of a start for you?

OpenStudy (anonymous):

@wolf1728 see that was the first thing I did, but the answer is 0.86, which again is not one of the options. That's why I'm totally confused.

OpenStudy (anonymous):

Perhaps explain to your teacher/tutor/mentor (whoever it may be) that you've done all necessary steps but cannot reach one of the supplied answers?

OpenStudy (wolf1728):

The answer is 0.8636757125 because 23^0.8636757125 = 15

OpenStudy (wolf1728):

Are you sure of the exact numbers in the problem? Could it be 2.3^x = 15 for example?